A standing wave has the following wave-function: y(x,t) = 0.2 sin(πx) cos⁡(12πt), where x and y are in meters, and t is in seconds. If the length of the string is L = 2 m and it is fixed at both ends, then the harmonic, n, in which the string is vibrating is:


n = 3


n = 5


n = 4


n = 6


n = 2


1
Expert's answer
2021-05-20T10:09:47-0400

"\frac{2\pi}{\lambda}=\pi\implies \lambda =2~m,"

"12\pi=2\pi\nu\implies \nu=6~Hz,"

"v=\lambda \nu=2\cdot 6=12~\frac ms,"

"\nu=\frac{nv}{2L}\iff 6=\frac{12n}{2\cdot 2}\implies \\n=2."